OpenStudy (anonymous):

What is the solution of the equation? -10x + sqrt(x+8) = -4

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ -10x + \sqrt{x+8} = -4 }\) \(\normalsize\color{blue}{ \sqrt{x+8} = 10x -4 }\) \(\normalsize\color{blue}{ x+8 = 100x^2 -80x+16 }\) \(\normalsize\color{blue}{ 0= 100x^2 -81x+8 }\) \(\normalsize\color{blue}{ 100x^2 -81x+8 =0 }\)

OpenStudy (solomonzelman):

use the quadratic formula from here